OpenStudy (anonymous):

Differentiate y=2^x sinx

OpenStudy (anonymous):

dy/dx = 2^x(cos x) + sin x(d/dx 2^x) = 2^x(cos x) + sin x(2^x ln 2) = 2^x(cos x + sin x ln 2) If you don't know how d/dx 2^x = 2^x ln 2, please say so.

OpenStudy (anonymous):

First step is using the product rule, if you don't notice.

OpenStudy (anonymous):

from the answer that i have it is actually we should rewrite it as \[\left| y \right|=2^x\left| sinx \right|\] why we need to put modulus/absolute sign for both sides?

OpenStudy (anonymous):

@PhoenixW, do you know why?

OpenStudy (anonymous):

well, actually i'm having the issue of when we need to use it as sometimes maybe we can know from any indicator from the questions itself.

OpenStudy (anonymous):

you get what i mean? it seems that from the answer we need to take modulus sign but if we were to answer this kind of questions, how do we know we need to use it?

OpenStudy (anonymous):

Honestly, the absolute value isn't required at all from what I can tell. I think I know why they assume so (regarding with integration of f'(x)/f(x)), but if I'm correct, they've assumed incorrectly.

OpenStudy (anonymous):

what do you mean by regarding with integration of f'(x)/f(x)?

OpenStudy (anonymous):

how are there two values of y for every value of x in y=sin x?

OpenStudy (anonymous):

you mean -1 and 1?

OpenStudy (anonymous):

I didn't think y=sin x had two values of y for each value of x, but PhoenixW just said there were. Is it something i'm not considering or remembering?

OpenStudy (anonymous):

complex numbers or something?

OpenStudy (anonymous):

i'm thinking if we were take sin (-1) it gives us negative value, therefore whenever we use natural logarithm of sin (-1) it is undefined. maybe that's why they put the modulus sign. am i correct?

OpenStudy (anonymous):

why would you have to take ln(sin x)?

OpenStudy (anonymous):

y=2^x sinx rewrite it as |y|=2x|sinx| then taking natural logarithm both sides and rearrange it gives you \[\ln \left| y \right|=x \ln 2 + \ln \left| \sin x \right|\]

OpenStudy (anonymous):

or you could use the product rule

OpenStudy (anonymous):

are you just trying to find the derivative or are you trying to find a differential equation?

OpenStudy (anonymous):

I can help you if it's just a derivative, but I don't know much about the concept of differential equations

OpenStudy (anonymous):

My bad on the two values of y, was thinking of a circle rather than just the sin function.

OpenStudy (anonymous):

If, for some reason, they use a method where they take ln of sin x before doing the product rule, they probably put the absolute value to prevent non-real numbers, in which case they'd need to absolute the y too.

OpenStudy (anonymous):

it's just the derivative of the equation. @Peter14, actually I don't understand why they need to put absolute sign for sinx and y? how would I know if i were to encounter the same question in future? is there any indicator for us to know that?

OpenStudy (anonymous):

@PhoenixW , i thinking like you too but still not sure..

OpenStudy (zzr0ck3r):

um y = sin(x) for sure is a function,

OpenStudy (zzr0ck3r):

im confused by the question you have ni21

OpenStudy (anonymous):

@zzr0ck3r, the question is actually differentiate y=2^x sinx. but i do have the answer. the only thing that i don't understand is why in the answer, the first step is to rewrite the question as |y|=2x|sinx|. then, from there it continues by taking natural logarithm both sides and rearranging it. it becomes as ln|y|=xln2+ln|sinx|. so again, my question is that why we need to use absolute value for both sides? why sinx?

OpenStudy (anonymous):

you don't need to use absolute value for both sides.

OpenStudy (anonymous):

hi @Peter14

OpenStudy (anonymous):

why do they have it in the solution? beats me.

OpenStudy (anonymous):

it might be to demonstrate a technique for dealing with products when you don't want to use product rule

OpenStudy (anonymous):

i'm not sure why... that's why i need your opinion

OpenStudy (anonymous):

that's the only thing I can think of. maybe ask your teacher?

OpenStudy (isaiah.feynman):